# Towards Trapped-Ion Thermometry Using Cavity-Based EIT

### v2

@@ V Determination of ion temperature and average phonon number
We studied this cavity-EIT phenomenon with the motivation of determining the ion temperature. It is evident from section IV that the linewidth of the cavity-EIT spectrum appears to serve as a probe parameter for temperature. The ion-temperature is incorporated in the systems by means of the thermal bath phonon occupancy ($n_{th}$) following Equation 7 and Equation 8 in steady state.
The simulation is performed for four different values of the thermal bath phonon occupancy, $n_{th}=0.5,1,5,$ and $10$, using the same set of parameters described in section III. A clear trend of increasing linewidth with increasing $n_{th}$ is observed (see 4(a)). This behavior indicates that the ion temperature can be mapped to the observed cavity-EIT linewidth. In other words, by analyzing the EIT width of the cavity-EIT spectrum, the ion temperature can be inferred.
In addition, the average phonon occupancy, $\bar{n}$, is an important quantity for various applications, such as estimating heating rates, determining fidelity of ion trap based quantum gates, indicating Lamb-Dicke regime etc. Since the phonon occupancy is directly related to temperature (see Equation 7), $\bar{n}$ can also be estimated from the cavity-EIT linewidth.
@@
Further, we demonstrate how the atom-cavity coupling strength affects the cavity-EIT linewidth, g, and Rabi frequency of the coupling laser. We provide linewidth maps in two dimensions for a range of mean thermal phonon number values. Appendix A contains the relevant plots.
VI Thermometry of a multi-ion system using a low finesse cavity
The discussion we have upto this point considers a single ion confined in a high-finesse cavity and operating in a strong coupling regime. However, operating in a strong-coupling regime is experimentally challenging with only a single ion, requiring either a near concentric-cavity regime [35] or fiber cavities where the polished surfaces are placed very close to each other [36]. Interestingly, both of these configurations possess very small mode volumes increasing the single atom cooperativity factor in a significant manner. Both are experimentally very challenging, with substantial efforts involved in making and aligning cavity mirrors with very high reflectivity.
@@ VI Thermometry of a multi-ion system using a low finesse cavity
Even in a weakly coupled cavity regime ($g\ll\kappa,\gamma$), it is possible to apply the proposed cavity EIT-based ion-thermometry method discussed earlier, albeit using more than one ion coupled to the cavity mode instead of a single one. The method works as one may achieve the collective strong coupling regime with a larger number of ions coupled to the cavity mode, thanks to the scaling of the single-atom cavity coupling factor ($g_{eff}\propto\sqrt{N}$) $g$ with ion number, $N$ [37]. Here g is the single-ion cavity coupling factor. The dependence of ion number on cavity-EIT linewidth is depicted in Figure 5. As the ion number increases, the cavity-EIT linewidth decreases exponentially, making the EIT linewidth even narrower than the single-atom case in the strong coupling regime. It suggests that using a sufficient number of atoms/ions collectively coupled to the cavity, one may be able to realize a very narrow cavity-EIT linewidth, which may be interesting for different other applications as well, such as using them as reference transitions in atomic clocks, laser frequency stabilization, etc.
@@
Figure 6: Temperature vs cavity-EIT linewidth for multi-atoms. The EIT linewidth is scaled with respect to the empty cavity width. The typical single-ion cavity EIT linewidth for one temperature is shown in the inset with the set of different parameters. Except temperature the same parameter has been chosen for the simulations performed in this plot.
@@
In the multi-atom/ion cavity system, the ion-cloud temperature may be estimated in a similar manner as outlined in the single ion cavity-EIT based thermometry scheme as discussed in section V. We have shown the variation in temperature as a function of cavity-EIT linewidth in fraction of empty cavity linewidth for different number of ions in Figure 6. The variation of the cavity EIT linewidth with temperature indicates that the linewidth is more sensitive to the temperature as one considers a higher number of atoms/ions coupled to the cavity mode. Thus, the proposed thermometry method is particularly well suited for implementation in low ion-cavity coupling strength cavity qed systems with multiple ions coupled to the cavity mode.
The current methodology discussed for the cavity EIT-based thermometry for a multi-ion cavity qed system has been developed with one simplifying assumption incorporated in our model. We assume that all the atoms/ions coupled to cavity mode possess the same thermal distribution or thermal occupancy. Modeling the system whereby different atoms/ions have a distinct thermal occupation is potentially challenging and computationally expensive. Additionally, it does not change the physics of the phenomena being modeled. Taking all ions to possess the same motional state is a thus a realistic assumption when all the ions are cooled to sufficiently low temperature ($\approx 10-50~\mu K$) and all of them share the same effective coupling to the cavity mode.
VII Roadmap towards experimental realization
@@
We present here an outline of experimentally implementing the trapped-ion thermometry method presented in this paper. As the thermal state detection method proposed is cavity EIT-based, one needs to engineer a trapped-ion cavity-QED system in the strong-coupling regime ($g^{2}\gg\kappa\gamma$). This may be achieved in two ways. The first approach would be to engineer a cavity configuration in the near-concentric or near-confocal regime as per the requirements of the physics objectives. We recommend using a high finesse (>50000>50000) cavity, such that the coupling factor ($g$) is high enough to justify operation in the strong coupling regime ($g^{2}\gg\kappa\gamma$) for a single atom/ion coupled to the cavity mode. The second approach focuses on a relatively relaxed regime of multi-ion based strong coupling regime of operation for a low finesse cavity.
@@
We need to choose a particular ion-species with suitable energy levels, and appropriate ion trap parameters such that the system stays in the resolved sideband regime ($\gamma\ll\omega_{sec}$). Most of the ion species used in cavity qed experiments globally typically have strong dipole-allowed transitions, with broad linewidths, where sidebands are not resolved in general. Such traps operate with typical trap frequencies in the range ($\omega_{sec}$) $\approx 0.5-2MHz$. So, one needs a judicious choice of the energy levels, such that the effective three-level system realized has very narrow linewidths for one of the excited levels.
@@
One may also consider using a dipole allowed transition (where typical linewidths are in the range of $\approx 5-20MHz$) for the trapped ion strongly coupled to the cavity field, one should consider operating the ion trap in a very tight confinement (very deep trap potential) of the ion with secular frequency greater transition linewidth ($\Gamma$). Such a mode of ion trap operation has been experimentally validated [38], where an 9Be+ ion was trapped with a very high RF voltage (600 V) and frequency (240 MHz), reaching secular frequencies in the range of $\sim 50~MHz$ and the sideband transitions were addressed in a dipole-allowed transition.
@@
If one considers applying the proposed thermometry method with multiple ions in low to medium finesse (finesse values in the range 3000150003000-15000) cavities having a relatively large mode waist ($50-100~\mu m$), about 10 ions coupled to the cavity mode will be sufficient to reach the strong coupling regime, assuming a single ion cavity coupling factor of $0.5$ MHz, cavity decay rate as well as the excited state decay rate about $1$ MHz.
VIII Conclusion
@@
We have presented a cavity-assisted thermometry scheme for trapped ions that takes advantage of the dependence of cavity-induced EIT on ions’ motional state. Through the explicit integration of thermalization and vibrational sidebands into a cavity-QED $\Lambda$-type system, we show that the mean phonon occupation and the ion temperature can be directly measured by the linewidth of the cavity-EIT transparency window. By means of motion-induced dephasing, thermal phonons perturb the coherence of the EIT dark state, resulting in a systematic expansion of the cavity-EIT resonance. This effect allows temperature readout directly from the cavity transmission spectrum, eliminating the need for projective internal-state detection in Rabi spectroscopy. Overall, for the diagnosis of ion temperature in cavity-QED systems, cavity-EIT-based thermometry provides a minimally invasive and experimentally feasible method. The established coupling between ion motion and cavity transmission goes beyond thermometry, offering a platform to investigate measurement-induced effects in cavity-QED systems and enabling continuous monitoring of motional heating.

### v3

@@ V Determination of ion temperature and average phonon number
We study the cavity-based EIT phenomenon with the motivation of determining the ion temperature. It is evident from section IV that the linewidth of the cavity-EIT spectrum appears to serve as a probe parameter for the trapped ion temperature. The ion temperature is incorporated in the systems by means of the thermal bath phonon occupancy ($n_{th}$) following Equation 4 and Equation 5 in steady-state.
The simulation is performed for four different values of the thermal bath phonon occupancy, $n_{th}=0.5,1,5,$ and $10$, using the same set of parameters described in section III. A clear trend of increasing linewidth with increasing $n_{th}$ is observed (see 4(a)). This behavior indicates that the ion temperature can be mapped to the observed cavity-EIT linewidth. In other words, by analyzing the EIT width of the cavity-EIT spectrum, the ion temperature can be inferred. In the subsequent sections, we discuss the applicability of this technique over a wide range of experimental parameters that govern current and futuristic cavity-QED systems in the strong and weak coupling regimes.
In addition, the average phonon occupancy, $\bar{n}$, is an important quantity for various applications, such as estimating heating rates, determining fidelity of ion trap-based quantum gates, indicating the Lamb-Dicke regime etc. Since the phonon occupancy is directly related to temperature (see Equation 4), $\bar{n}$ can also be estimated from the cavity-EIT linewidth.
Figure 4: (a) Cavity-EIT spectrum for different $n_{th}$ value (b) Ion temperature and average phonon occupancy ($\bar{n}$) as a function of cavity-EIT linewidth. The cavity decay rate, $\kappa=0.4MHz$, cavity coupling factor, $g=3\kappa$. The decay rates from the excited states are taken as $\gamma_{eg}=\gamma_{eu}=\kappa$. The phonon coupling rate is taken as $\gamma_{b}=0.25\kappa$.
@@
Further, we demonstrate how the ion-cavity coupling strength affects the cavity-EIT linewidth, $g$, and Rabi frequency of the coupling laser. We provide linewidth maps in two dimensions for a range of mean thermal phonon number values. Appendix C contains the relevant discussion and the requisite plots.
VI Thermometry of a multi-ion cavity-QED system in the weak coupling regime
The discussion so far has focused on a single ion confined in a high-finesse cavity and operating in the strong coupling regime. However, operating in the strong-coupling regime is experimentally challenging with only a single ion, requiring either a near concentric-cavity regime [42] or fiber cavities where the polished surfaces are placed very close to each other [43]. Interestingly, both of these configurations possess very small mode volumes leading to an increase in the single-atom cooperativity factor in a significant manner. Both regimes of operation are experimentally challenging, with substantial effort necessary in making and aligning cavity mirrors with very high reflectivity for the above two cavity configurations.
@@ VI Thermometry of a multi-ion cavity-QED system in the weak coupling regime
Our analysis reveals that it is possible to perform ion thermometry measurements for multi-ion cavity-QED systems having a relatively low single-ion cavity coupling factor ($g\ll\kappa,\gamma$) that places such systems in the weak coupling regime ($g\ll\kappa,\gamma$) for a single ion. We show through numerical simulations that is possible to apply the proposed cavity EIT-based ion-thermometry method discussed earlier even for such systems in the weak coupling regime, albeit using more than one ion coupled to the cavity mode instead of a single one. Using a multi-ion configuration coupled to the cavity mode places the system effectively in the strong coupling regime due to the collective ($g_{eff}\propto\sqrt{N}$) scaling up of the single-ion cavity coupling factor.
The Hamiltonian for a multi-ion cavity-QED system may be stated as
$H=$ $\Delta_{p}S_{gg}-\Delta_{p}a^{\dagger}a+g\left(a^{\dagger}S_{ge}+aS_{eg}\right)$ (10)
$+\eta\Omega_{c}\left(S_{eu}b^{\dagger}+S_{ue}b\right)+\epsilon\left(a^{\dagger}+a\right).$
where,
$S_{ij}=\sum_{k=1}^{N_{\text{ion}}}\left|i\right\rangle^{(k)}\left\langle j\right|^{(k)}=\sum_{k=1}^{N_{\text{ion}}}\sigma_{ij}^{(k)}.$
are the collective raising and lowering ionic operators for $i\neq j$, and ionic energy-level population operators for $i=j$.
The system dynamics, including the decoherence and dissipative effects, are determined by the master equation as,
$\frac{d\rho}{dt}=$ $-i[H,\rho]+\kappa\left(2a\rho a^{\dagger}-a^{\dagger}a\rho-\rho a^{\dagger}a\right)$ (11)
$+\gamma_{eu}\sum_{k=1}^{N_{\text{ion}}}\left(2\sigma_{ue}^{(k)}\rho\sigma_{eu}^{(k)}-\sigma_{ee}^{(k)}\rho-\rho\sigma_{ee}^{(k)}\right)$
$+\gamma_{eg}\sum_{k=1}^{N_{\text{ion}}}\left(2\sigma_{ge}^{(k)}\rho\sigma_{eg}^{(k)}-\sigma_{ee}^{(k)}\rho-\rho\sigma_{ee}^{(k)}\right)$
$+\gamma_{u,deph}\sum_{k=1}^{N_{\text{ion}}}\left(2\sigma_{uu}^{(k)}\rho\sigma_{uu}^{(k)}-\sigma_{uu}^{(k)}\rho-\rho\sigma_{uu}^{(k)}\right)$
$+\gamma_{b}(n_{\mathrm{th}}+1)\left(2b\rho b^{\dagger}-b^{\dagger}b\rho-\rho b^{\dagger}b\right)$
$+\gamma_{b}n_{\mathrm{th}}\left(2b^{\dagger}\rho b-bb^{\dagger}\rho-\rho bb^{\dagger}\right).$
@@
The method works as one may achieve the collective strong coupling regime with a larger number of ions coupled to the cavity mode, thanks to the scaling of the single-atom cavity coupling factor ($g_{eff}\propto\sqrt{N}$) $g$ with ion number, $N$ [44]. Here $g$ is the single-ion cavity coupling factor. The dependence of ion number on cavity-EIT linewidth is depicted in Figure 5. As the ion number increases, the cavity-EIT linewidth decreases exponentially, making the EIT linewidth even narrower than the single-atom case in the strong coupling regime. It suggests that using a sufficient number of atoms/ions collectively coupled to the cavity, one may be able to realize a very narrow cavity-EIT linewidth, which might be useful for a variety of other applications as well, such as using them as reference transitions in atomic clocks, laser frequency stabilization, etc.
@@
Figure 6: Temperature vs cavity-EIT linewidth for multi-ions. The EIT linewidth is scaled with respect to the empty cavity width. The typical single-ion cavity EIT linewidth for one temperature is shown in the inset with the set of different parameters. Except for temperature, the same parameters have been chosen for the simulations performed in this plot.
@@
In the multi-atom/ion cavity system, the ion-cloud temperature may be estimated in a similar manner as outlined in the single-ion cavity EIT-based thermometry scheme as discussed in section V. We have shown the variation in temperature as a function of cavity-EIT linewidth in a fraction of empty cavity linewidth for different numbers of ions in Figure 6. The variation of the cavity EIT linewidth with temperature indicates that the linewidth is more sensitive to the temperature as one considers a higher number of atoms/ions coupled to the cavity mode. Thus, the proposed thermometry method is particularly well-suited for cavity-QED systems with low ion-cavity coupling strength, especially when multiple ions are coupled to the cavity mode.
In the multi-ion model, we assume that all ions share an identical thermal occupation. This approximation is justified under a uniformly cooled condition, where all the ions are cooled to a sufficiently low temperature ($\approx 10-50\mu K$) and undergo comparable coupling to the cavity field. Consequently, the collective cavity response can be described using a thermal distribution reminiscent of a single ion.
VII Thermometry of a multi-ion cavity-QED system with a large excited-state decay rate
The three-level system we have considered for our ion thermometry is an ideal system, with a very low decay rate from the excited state. However, most ion species used in trapped-ion experiments have a broad excited-state linewidth (i.e., a large decay rate), which helps cool the ions efficiently with Doppler cooling.
Figure 7: Cavity-EIT linewidth variation with the number of ions for the case of high decay rate. The simulation parameters are shown in the inset.
A large decay rate induces decoherence in the system, which washes out the cavity-EIT effect, and only the bare cavity signal remains visible at the output. Several groups worldwide have demonstrated the cavity-EIT effect using ion/atom species with a large decay rate from the excited state [44, 19]. The experimental results from these systems show that for a single ion the cavity-EIT signal is not clearly visible; however, increasing the number of atoms leads to the appearance of cavity-EIT peaks. This occurs because, although the ion-cavity system is not configured in the strong coupling regime for a single ion/atom (as the decay rates exceed the cavity coupling factor), the effective coupling strength increases with multiple ions coupled to the intracavity field. Consequently, the collective ion-cavity coupling can overcome the large decay rates of the excited states, as discussed in section VI. More discussion about their system and an excellent agreement of our simulation results with the analytical expression and the reported experimental data can be found in Appendix A and B.
Motivated by these studies, we extend our ion-thermometry method to cases where the excited state decays rapidly to the ground states in a three-level lambda system. The cavity-EIT peaks become visible when considering a multi-ion system, and the linewidth decreases with the increasing number of atoms, which is the usual behavior similar to that obtained in section VI. The variation of the cavity linewidth with the number of atoms, obtained from simulations performed for up to 500 ions, along with the corresponding simulation parameters, is depicted in Figure 7.
[…]
Interestingly, when running the simulations at different temperatures, i.e., for different values of $n_{th}$, the linewidth does not vary at all. This occurs because, for systems with large decay rates, the resolved sideband condition ($\gamma\ll\omega_{sec}$) is no longer satisfied. As a result, the motional sidebands cannot be resolved, which is essential for the thermometry process. Therefore, we conclude that, in addition to strong coupling, a necessary condition for implementing this thermometry method is that the ion must operate in the resolved sideband regime. The only remaining option in such cases is to increase the trap frequency $\omega_{c}$; in other words, a tighter confinement of the ion with a deeper trapping potential is required.
To assess the applicability of the thermometry method in the presence of strong excited-state decay, we analyze the temperature dependence of the cavity-EIT linewidth for different numbers of ions as shown in Figure 8. A distinct and monotonic dependence of the normalized linewidth on the temperature is preserved even for large decay rates ($\gamma_{eu}=20\kappa$, $\gamma_{eg}=7\kappa$). Although the decoherence caused by spontaneous emission, which also partially broadens the EIT feature, collective enhancement ($g_{eff}\propto\sqrt{N}$) strengthens the effective light-matter interaction to compensate for decoherence. These results demonstrate that the suggested cavity-EIT thermometry method is robust even in regimes of significant radiative decay.
VIII Roadmap towards experimental realization
@@
We discuss an outline of experimental implementations for the trapped-ion thermometry method discussed in the previous sections of the article. As the thermal state detection method proposed is cavity EIT-based, one needs to engineer a trapped-ion cavity-QED system in the strong-coupling regime ($g^{2}\gg\kappa\gamma$). This may be achieved in two ways. The first approach would be to engineer a cavity configuration in the near-concentric or near-confocal regime as per the requirements of the physics objectives. We recommend using a high finesse (>50000>50000) cavity, such that the coupling factor ($g$) is high enough to justify operation in the strong coupling regime ($g^{2}\gg\kappa\gamma$) for a single atom/ion coupled to the cavity mode. The second approach focuses on a relatively relaxed regime of multi-ion based strong coupling regime of operation for a low-finesse cavity.
@@
We need to choose a particular ion species with suitable energy levels, and appropriate ion trap parameters such that the system stays in the resolved sideband regime ($\gamma\ll\omega_{sec}$). Most of the ion species used in cavity-QED experiments globally typically have strong dipole-allowed transitions, with broad linewidths, where sidebands are not resolved in general. Such traps operate with typical trap frequencies in the range ($\omega_{sec}$) $\approx 0.5-2MHz$. So, one needs a judicious choice of the energy levels, such that the effective three-level system realized has very narrow linewidths for one of the excited levels.
@@
One may also consider using a dipole allowed transition (where typical linewidths are in the range of $\approx 5-20MHz$) for the trapped ion strongly coupled to the cavity field, one should consider operating the ion trap in a very tight confinement (very deep trap potential) of the ion with secular frequency greater than the transition linewidth ($\Gamma$). Such a mode of ion trap operation has been experimentally validated [45], where an 9Be+ ion was trapped with a very high RF voltage (600 V) and frequency (240 MHz), reaching secular frequencies in the range of $\sim 50~MHz$ and the sideband transitions were addressed in a dipole-allowed transition.
@@
If one considers applying the proposed thermometry method with multiple ions in low- to medium-finesse (finesse values in the range 3000150003000-15000) cavities having a relatively large mode waist ($50-100~\mu m$), about 10 ions coupled to the cavity mode can allow the system to operate in the strong-coupling regime, assuming a single ion cavity coupling factor of $0.5$ MHz and cavity decay rate as well as the excited state decay rate of about $1$ MHz.
IX Conclusion
@@
We have presented a cavity-assisted thermometry scheme for trapped ions that takes advantage of the dependence of cavity-induced EIT on ions’ motional state. Through the explicit integration of thermalization and vibrational sidebands into a cavity-QED $\Lambda$-type system, we show that the mean phonon occupation and the ion temperature can be directly measured by the linewidth of the cavity-EIT transparency window. By means of motion-induced dephasing, thermal phonons perturb the coherence of the EIT dark state, resulting in a systematic expansion of the cavity-EIT resonance. This enables direct extraction of the ion temperature from the measured cavity transmission spectrum, eliminating the need for projective internal-state detection in Rabi spectroscopy. Overall, for the diagnosis of ion temperature in cavity-QED systems, cavity-EIT-based thermometry provides a minimally invasive and experimentally feasible method. The established coupling between ion motion and cavity transmission goes beyond thermometry, offering a platform to investigate measurement-induced effects in cavity-QED systems and enabling continuous monitoring of motional heating
@@ Acknowledgments
Appendix A Transmission and reflection spectra with a bare cavity
Figure 9: Cavity reflection from simulation, analytical expression and published experimental data [46] as a function of probe detuning. The free spectral range of the cavity is $\nu_{FSR}=12.7~\text{GHz}$. The transmission coefficients of the mirrors are $T_{1}=1500~\text{ppm}$ and $T_{2}=5~\text{ppm}$, while the intracavity loss coefficient is $A=600~\text{ppm}$. For these parameters, the resulting cavity finesse is approximately $F\sim 3000$.
To validate the theoretical model and numerical simulations discussed in the earlier sections of the article, we examine the empty-cavity limit by setting the ion-cavity coupling strength to zero and ignoring all phonon-related contributions in the Hamiltonian. Under these conditions, the system reduces to a driven, dissipative optical cavity described solely by the bare cavity Hamiltonian and the decay channels associated with it.
The Hamiltonian of the bare cavity system can be expressed as,
$H_{\text{bare}}=\Delta_{c}a^{\dagger}a+i\sqrt{2\kappa_{1}}(a^{\dagger}a_{in}-a_{in}^{\dagger}a),$ (12)
where $a^{\dagger}$ ($a$) is the creation (annihilation) operator for the cavity mode, $\Delta_{c}$ represents the probe-cavity detuning, and $a_{in}$ is the input field injected into the cavity.
Standard input-output relations are used to calculate the resulting transmission ($\mathcal{T}$) and reflection ($\mathcal{R}$) spectra. For an empty cavity, the response follows the well-known Lorentzian resonance profile determined by the cavity decay rates.
$\mathcal{R}=\left|\frac{a_{\mathrm{ref}}}{a_{\mathrm{in}}}\right|^{2}=\frac{(\kappa-2\kappa_{1})^{2}+\Delta_{c}^{2}}{\kappa^{2}+\Delta_{c}^{2}},$ (13)
$\mathcal{T}=\left|\frac{a_{\mathrm{trans}}}{a_{\mathrm{in}}}\right|^{2}=\frac{4\kappa_{1}\kappa_{2}}{\kappa^{2}+\Delta_{c}^{2}},$ (14)
where, $\kappa$ corresponds to the total decay rate of the cavity field, while $\kappa_{1}$ and $\kappa_{2}$ represent the contribution to the cavity loss arising from the imperfect reflectivity of the two cavity mirrors.
Figure 9 shows the cavity transmission and reflection spectra obtained from our numerical simulation together with the corresponding analytical expression and previously reported experimental data [46]. The obtained spectra exhibit the expected Lorentzian lineshape with normalization, reproducing the well-known analytical response of a bare cavity. The validity of the reduced Hamiltonian reflecting the response of the bare cavity and the numerical method used throughout the manuscript are confirmed by the excellent agreement with the standard bare-cavity reflection and transmission profiles. The bare cavity response therefore serves as a baseline reference against which the modifications arising from ion-cavity interactions and phonon-induced effects can be reliably identified in the main results.
Appendix B Comparison of simulation results with analytical expressions and experimental data
We verify our numerical simulations for a multi-ion cavity-QED system exhibiting a low single-ion cavity coupling factor as well as using ions with a large decay rate of the excited state. These systems can be driven in the collective strong coupling regime by appropriate choice of ions coupled to the cavity mode. We compare the simulated normalized cavity emission spectrum with previously published experimental data and results and compare them with available analytical expressions mentioned in [41, 46].
[…]
$H=\hbar\Big[$ $\Delta_{1}\sigma_{ee}+(\Delta_{1}-\Delta_{2})\sigma_{uu}+g(a^{\dagger}\sigma_{ge}+\sigma_{eg}a)$ (15)
$+\frac{\Omega_{c}}{2}(\sigma_{ue}+\sigma_{eu})+\epsilon(a^{\dagger}+a)-\Delta a^{\dagger}a\Big],$
where detunings are defined as $\Delta_{1}=\omega_{eg}-\omega_{p}$, $\Delta_{2}=\omega_{eu}-\omega_{c}$ and $\Delta=\omega_{c}-\omega_{p}$. $\omega_{p}$ and $\omega_{c}$ correspond to the probe and control field frequencies, respectively.
The intracavity photon number can be obtained analytically from the steady-state cavity field in the weak-probe limit as,
$\langle n\rangle=\langle a^{\dagger}a\rangle=\frac{|\epsilon|^{2}}{\left|(\Delta+i\kappa)-\chi\right|^{2}},$ (16)
where $\kappa$ represents the cavity decay rate and $\chi$ is the effective susceptibility of the EIT medium,
$\chi=\frac{g^{2}N(\Delta-\Delta_{1}+\Delta_{2}+i\gamma_{u,\mathrm{deph}})}{(\Delta-\Delta_{1}+i(\gamma_{eg}+\gamma_{eu}))}$ (17)
$\times\frac{1}{(\Delta-\Delta_{1}+\Delta_{2}+i\gamma_{u,\mathrm{deph}})-\Omega_{c}^{2}/4}.$
With these analytical expression and the parameters listed in [41] we compute the transmission spectrum and compared it with the outcomes of our numerical model. Figure 10 shows the normalized cavity emission as a function of probe detuning $(\Delta/\kappa)$ for $N_{ions}=17$. As observed, the simulation accurately reproduces the key spectral features, including the central transparency peak and two sidebands.
Figure 10: Normalized cavity emission from simulation, analytical expression and published experimental data [41] as a function of probe detuning. The cavity decay rate ($\kappa$), is 1 MHz, while the single-atom cavity coupling rate ($g_{0}$) is $0.7\kappa$. The control field Rabi frequency $(\Omega_{c})$ is $3MHz$. The detuning for probe laser ($\Delta_{1}$) and control laser ($\Delta_{2}$) are $0.8\kappa$ and $0.55\kappa$ respectively.
The accuracy of the theoretical implementation is confirmed by the excellent agreement between the simulations and the analytical prediction. Furthermore, the close overlap with experimental data confirms the physical assumptions incorporated in the model.This agreement shows that the numerical framework reproduces experimentally observed behavior with high reliability and captures the cavity-EIT dynamics.
Figure 11: Cavity reflection from simulation, analytical expression and published experimental data [46] as a function of probe detuning. The decay rate for 1st mirror ($\kappa_{1}$), 2nd mirror ($\kappa_{2}$) are $1.6MHz$ and $0.5kHz$ respectively. The decay rates from excited state to first and 2nd ground state are, $\gamma_{31}=12.6~MHz$ and $\gamma_{21}=0.6~kHz$ respectively. The single atom coupling factor g=0.48MHzg=0.48~MHz, while control field Rabi frequency, $\Omega_{c}=4.1~MHz$.
To further validate our numerical model, we also compare our results with the reflection spectra published in [46]. This work examines the cavity reflection from an ensemble of atoms coupled to the cavity mode in contrast to the previous work in [41], which concentrates on the cavity transmission spectrum. Specifically, for the system of about 500 atoms, the reflection spectrum was investigated analytically and experimentally in [46].
The cavity-EIT spectrum shown in Figure 11 is deduced from the analytical expression of the normalized cavity reflectivity with respect to the input field, as explained in [46]. The reflectivity can be written as,
$R$ $=\left|{\frac{a_{ref}}{a_{in}}}\right|^{2}$ (18)
$=\frac{\left[\kappa_{1}-\kappa_{2}-\kappa_{A}-\mathrm{Im}(\chi_{\Lambda})\right]^{2}+\left[\Delta-\mathrm{Re}(\chi_{\Lambda})\right]^{2}}{\kappa^{2}+\Delta^{2}+\left|\chi_{\Lambda}\right|^{2}+2\left[\kappa\,\mathrm{Im}(\chi_{\Lambda})-\Delta\,\mathrm{Re}(\chi_{\Lambda})\right]}$
Here $a_{ref}$ and $a_{in}$ are the reflected and input field amplitude respectively. $\kappa_{1}$, $\kappa_{2}$ and $\kappa_{A}$ are the decay rates for field losses from the cavity first mirror, second mirror and additional internal losses due to various mechanisms. $\Delta$ is the probe field detuning and $\kappa$ is the total decay rate ($\kappa=\kappa_{1}+\kappa_{2}+\kappa_{A}$). $\chi_{\Lambda}$ is the three-level non-linear susceptibility of the atomic ensemble, which can be expressed as,
$\chi_{\Lambda}=\frac{ig^{2}N}{\gamma+i\Delta}\frac{\ln(1+s)}{s}$ (19)
Where, $g$ is the single ion coupling factor, $N$ is the effective number of atoms coupled with the cavity field, $\gamma$ is the combined decay rates for the excited state. The saturation parameter for the two-photon transition is $s$, which can be articulated as,
$s=\frac{\Omega_{C}^{2}/2}{(\gamma_{12}+i\delta)(\gamma+i\Delta)}$ (20)
Where, $\Omega_{c}$, is the control field Rabi frequency, $\gamma_{12}$ is the dephasing rate between the ground states. $\delta$ is the two-photon detuning, which the difference between probe laser detuning ($\Delta$) and control laser detuning ($\Delta_{32}$), $\delta=\Delta-\Delta_{32}$.